OpenStudy (anonymous):

Factor x^2 + 64. Check your work. Factor 16x^2 + 49. Check your work. Find the product of (x + 9i)^2 Find the product of (x – 2i)^2 Find the product of (x + (3+5i))^2 Please help asap!

OpenStudy (anonymous):

\[x^2+64=x^2-(-64)=x^2-64\iota^2=x^2-\left( 8\iota \right)^2=\left( x+8\iota \right)\left( x-8\iota \right)\]

OpenStudy (anonymous):

is that x^2+64 completely factored?

OpenStudy (anonymous):

Are you sure it's not 16^2 - 49?

OpenStudy (anonymous):

(x + 9i)^2 = x^2 + 18ix + 81i